OpenStudy (tumblewolf):

I need help with using limit definition to find the tangent line. Will fan and medal! I did it in derivative form and my teacher wants it the other way but I can't figure it out. F(x)=2x^-6x+2 and x=2 And don't know where to go from here. Thank you!! ((2h2+8h+8−12+6h+2)−8−12+2)/h this is what I have so far

myininaya (myininaya):

derivative form?

myininaya (myininaya):

do you mean by using the definition of derivative?

OpenStudy (tumblewolf):

That's just what she called it. She said I did it in the wrong form

myininaya (myininaya):

I think you mean by definition of derivative derivative form means nothing to me

myininaya (myininaya):

Also I think there is a type-o in your function

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0} \frac{f(x+h)-f(x)}{h}\] this is definition of derivative

OpenStudy (tumblewolf):

Yes

myininaya (myininaya):

and she probably means she wants you to find the derivative using this definition

OpenStudy (tumblewolf):

Yes

myininaya (myininaya):

\[f(x)=2x^{-6x}+2\] is this really the correct function?

OpenStudy (tumblewolf):

\[2x^2-6+2\]

myininaya (myininaya):

that is weird why wouldn't they just combine like terms

OpenStudy (tumblewolf):

\[2x^2-6x+2\]

myininaya (myininaya):

\[f(x)=2x^2-6x+2 \\ f(x+h)=2(x+h)^2-6(x+h)+2\] just should have been the first thing stated

myininaya (myininaya):

\[\frac{f(x+h)-f(x)}{h} \\ \frac{[2(x+h)^2-6(x+h)+2]-[2x^2-6x+2]}{h} \\ \text{ you can do some multiplying out and canceling of terms in numerator first } \\ \text{ I sometimes like to factor things by rearranging the right terms together }\]

OpenStudy (tumblewolf):

So combine the x's?

myininaya (myininaya):

if you choose factoring way.. combine terms like this first: \[\frac{2[(x+h)^2-x^2]-6[(x+h)-x]+[2-2]}{h}\] if you choose to multiply things out... you should have this next: \[\frac{2x^2+4xh+2h^2-6x-6h+2-2x^2+6x-2}{h} \]

myininaya (myininaya):

and then for the multiplying things out part you collect like terms in the numerator

myininaya (myininaya):

for the factoring way if you choose it you will need to know how to factor a difference of squares

OpenStudy (tumblewolf):

I am still really confused on this

myininaya (myininaya):

on multiplying part?

OpenStudy (tumblewolf):

On all of it and how things plug in

myininaya (myininaya):

\[f(x+h)=2(x+h)^2-6(x+h)+2 \\ f(x+h)=2(x+h)(x+h)-6(x+h)+2\] do you know know to expand (x+h)(x+h) through multiplication and -6(x+h) through multiplication ?

myininaya (myininaya):

oh you aren't sure how to find f(x+h)?

OpenStudy (tumblewolf):

I think it's the equation but I'm not sure

myininaya (myininaya):

\[f(x)=2x^2-6x+2\] notice the x's.. the x acts as the input value until we can think of another input value to replace it with

OpenStudy (tumblewolf):

\[(2h^2+8h+8-12+6h+2)-8-12+2)/h\]

myininaya (myininaya):

\[f(x)=2x^2-6x+2 \\ \text{ examples of how the machine's outputs change base on the input } \\ f(1)=2(1)^2-6(1)+2 \\ f(1+2)=2(1+2)^2-6(1+2)+2 \\ f(3)=2(3)^2-6(3)+2 \\ f(\star)=2(\star)^2-6(\star)+2 \\ f(\star+\Delta)=2(\star+\Delta)^2-6(\star+\Delta)+2 \\ f(x+h)=2(x+h)^2-6(x+h)+2\] we are finding f(x+h) since in the definition of the derivative it has f(x+h) ... \[f'(x)=\lim_{h \rightarrow 0} \frac{\color{red}{f(x+h)}-f(x)}{h}\]

myininaya (myininaya):

\[\text{ everywhere there is an } x \text{ in } f(x)=2x^2-6x+2 \\ \text{ we replace it with } (x+h) \text{ since we are to plug in } f(x+h)\]

OpenStudy (tumblewolf):

I get that part

myininaya (myininaya):

\[f(x+h)=2(x+h)^2-6(x+h)+2\] can you expand this (perform the multiplication)

OpenStudy (tumblewolf):

2x+2h+2x+2h-6x+6h+2

myininaya (myininaya):

how did you get that?

OpenStudy (tumblewolf):

Just multiplying it out

myininaya (myininaya):

how did you multiply 2(x+h)^2 out?

OpenStudy (tumblewolf):

2(x+h)(x+h) Then just kinda foiled it

myininaya (myininaya):

but your x^2, h^2, and 2xh term disappeared? 2(x+h)(x+h) 2(x^2+2xh+h^2)

OpenStudy (tumblewolf):

oh

myininaya (myininaya):

do you understand how to multiply binomials?

myininaya (myininaya):

\[(a+b)(x+y) \\ a(x+y)+b(x+y) \\ ax+ay+bx+by \\ \]

OpenStudy (tumblewolf):

yes

myininaya (myininaya):

\[(x+h)(x+h) \\ x(x+h)+h(x+h) \\ xx+xh+hx+hh \\ x^2+2xh+h^2\]

OpenStudy (tumblewolf):

yes

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0} \frac{[f(x+h)]-[f(x)]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{[2(x^2+2xh+h^2)-6x-6h+2]-[2x^2-6x+2]}{h}\] do distributive property and then combine like terms

OpenStudy (tumblewolf):

\[(2x^2+4xh+2h^2)-6x-6h+2)-2x^2-6x+2\]

myininaya (myininaya):

the whole reason I put [ ] around the 2x^2-6x+2 was so you wouldn't forget that the whole f(x) is being subtract not just part of

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0} \frac{[f(x+h)]-[f(x)]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{[2(x^2+2xh+h^2)-6x-6h+2]-[2x^2-6x+2]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{2x^2+4xh+2h^2-6x-6h+2-2x^2+6x-2}{h}\]

myininaya (myininaya):

combine like terms in numerator

OpenStudy (tumblewolf):

\[(4xh+6h)/h\]

myininaya (myininaya):

missing one more term

myininaya (myininaya):

and also there is a - in front of 6h

OpenStudy (tumblewolf):

\[(4xh+2h^2-6h)/h\]

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0} \frac{[f(x+h)]-[f(x)]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{[2(x^2+2xh+h^2)-6x-6h+2]-[2x^2-6x+2]}{h} \\ f'(x)=\lim_{h \rightarrow 0} \frac{\color{Red}{2x^2}+4xh+2h^2-\color{blue}{6x}-6h+\color{green}{2}-\color{red}{2x^2}+\color{blue}{6x}-\color{green}{2}}{h} \\ \] correct

myininaya (myininaya):

now you should see a common factor from numerator and denominator

myininaya (myininaya):

h/h=1

myininaya (myininaya):

\[\frac{h}{h} \cdot \frac{4x+4h-6}{1}\]

OpenStudy (tumblewolf):

Doesn't h=0 though?

myininaya (myininaya):

we will plug in h=0 once we get rid of the thing h on bottom

myininaya (myininaya):

0/0 is indeterminate this is why getting rid of that h in the denominator will give us a form that is non-indeterminate

OpenStudy (tumblewolf):

4x-6?

OpenStudy (tumblewolf):

4x+4-6

myininaya (myininaya):

are you understanding we have 4x-6?

OpenStudy (tumblewolf):

I was canceling out the h

OpenStudy (tumblewolf):

So that gives us 6?

OpenStudy (tumblewolf):

Is it just y=4x-6?

OpenStudy (tumblewolf):

@myininaya

myininaya (myininaya):

\[\frac{4xh+2h^2-6h}{h}=\frac{h}{h} \cdot \frac{4x+2h-6}{1}=(1) \cdot (4x+2h-6) =4x-6 \text{ since } h \rightarrow 0\]

myininaya (myininaya):

so yes this is how I got f'(x) above

myininaya (myininaya):

f'(x)=4x-6

myininaya (myininaya):

f'(x)=4x-6 is the slope formula

myininaya (myininaya):

function

myininaya (myininaya):

so if we want the slope at x=5 evaluate f'(5)

OpenStudy (tumblewolf):

Thank you